One positive number is 5 times another number. The difference between the two numbers is 356, find the numbers
Answer by stanbon(75887)

One positive number is 5 times another number. The difference between the two numbers is 356, find the numbers
--------
Equations:
x = 5y
x - y = 356
---------
Substitute for "x" and solve for "y":
5y-y = 356
4y = 356
y = 89
----
Solve for "x":
x = 5*89 = 445
======================
Cheers,
Stan H.